# A Survey on Dynamic Neural Networks: from Computer Vision to Multi-modal Sensor Fusion

### v1

@@ 2.2.2 Mixture of Experts
Wang et al. (2020) presents an architecture based on MoE that composes the filter of a CNN on the fly from a set of (expert) channels. The input is fed into both a base Convolutional Network and a shallow embedding network. The embedding network outputs a latent mixture of weights, which is used by the convolutional network to select the expert channels to use for each specific layer. The paper also demonstrates that this gating method is equivalent to constructing a MoE for each output channel. A loss based on the computational cost is also used to encourage sparsity in the gating network. Also, Hua et al. (2019) presents a similar idea and introduces a revision of the MoE method to learn specialized convolutional kernels for each example. This increases the size and capacity of the network while maintaining efficient inference. The idea is to compute convolutional kernels as a linear combination of $n$ learned experts. In this way, the convolutional kernel is constructed on the fly based on the input needs.
@@
In Hazimeh et al. (2021) the authors raise the problem of slow convergence in the case of sparse gates (like top-$k$), which results in performance issues when training with gradient-based methods. The authors then introduce a reformulation of the binary selection based on first-order methods (such as SGD). Zhang et al. (2023) presents a method to strengthen the MoE CNNs. The paper talks mostly about adversarial attacks and robustness, trying to understand what, among all the factors (network, router), is mostly responsible for the lack of robustness and proposing a method to train robust MoE CNNs.
Riquelme et al. (2021) is the first to introduce the use of MoE in the context of ViT models. The specialization in this case can be performed on a token level, where each token gets assigned to one (or more) experts, and experts can accept tokens up to a capacity level. Generally, in ViTs, the experts is applied to the MLP portion of the network. The paper also introduces a routing algorithm that repurposes model sparsity to skip the computation of some patches, reducing computations on uninformative image regions. Liu et al. (2024) further studies different routing approaches of MoE in the context of ViTs. The comparison considers binary routers for hard assignment among experts, and soft routers for soft assignment between experts and weighted combinations of the expert output tokens. Further assignment divisions can be made, either matching experts to tokens or vice versa. The study arrives at the conclusion that the soft MoE outperforms the binary solution in terms of accuracy and thus is favorable.
Last, M3ViT (Liang et al., 2022) introduces the use of a MoE framework to accelerate multi-task ViTs. Sparse task-specific experts get activated during training, which disentangles the parameter spaces to avoid different tasks’ training conflicts. The solution proposes tokens to per-expert queues to enable expert-by-expert computation rather than token-by-token. At inference time, this allows for activating only the task-corresponding sparse expert. The solution proposes also a computation reordering scheme to switch between tasks for deployment on memory-constrained devices.
@@ 2.2.3 Applications
In this section, we go through publications that present the application of computational routing techniques to solve specific applicational problems. Park et al. (2021) explore the use of dynamic routing networks in the context of object detection. The proposed architecture is based on MSDNet (Huang et al., 2018) and automatically generates a sample-adaptive routing path on the fly. Two additional metrics for a global budget constraint and local path similarity regularization are added. The first serves the purpose of inducing sparsity in the network, while the second aims to generate more diverse routing paths.
@@
Pavlitskaya et al. (2020) investigates a MoE architecture to achieve interpretability in the context of semantic segmentation while retaining comparable result quality. The principle consists in being able to observe both the overall model output as well as individual expert outputs, where the agreement or disagreement between those partial solutions can be used for insights into the decision process.
@@
In Sun et al. (2021), a framework for efficient video recognition is presented. The solution selects the optimal precision for each frame conditioned on the input. Through Gumbel-Softmax sampling, an end-to-end differentiable approach selects optimal precision, taking both accuracy and efficiency into account. The solution is composed of a lightweight policy network (composed of a feature extractor and a LSTM module) and a video recognition network. The method also allows to skip frames in addition to selecting the precision dynamically. Furthermore, knowledge distillation from the full-precision recognition network is performed to guide the training of lower precisions and to improve accuracy. Yang et al. (2024) presents DyFADet, a module that can adapt both kernel weights and receptive fields at different timestamps in the task of temporal action detection. The architectures dynamically adjust detector parameters when applied at different pyramid levels, which corresponds to detecting the actions with different time ranges.
@@
Furthermore, Wang et al. (2021) introduces an adaptive patch localization solution for video recognition. By using a Reinforcement Learning approach, the patch localization problem is defined as a sequential decision task and a CNN is used to quickly process the full video sequence. Then, the selected patches are used by a larger network to get the final prediction. The solution takes into account also the possibility of skipping some less valuable frames in the training policy. Wang et al. (2022) reformulates AdaFocus as one-stage algorithm by introducing a differentiable interpolation-based patch selection approach. Furthermore, in this case, a conditional-exit technique is included to perform temporal adaptive computation. Wang et al. (2022) improves on top of the second version of AdaFocus by activating the expensive high-capacity network only on some small but informative 3D video cubes.
@@
Paper Summary Architecture Dynamic component Section
Ahmed et al. (2016) Introduces MoE architecture for CNN CNN MoE 2.2.2
Bolukbasi et al. (2017) Ensemble of networks chained in an acyclic computation graph CNN Computational path 2.2.1
Figurnov et al. (2017) Dynamically adjusts the number of layers for certain regions of the image CNN Layer 2.2.1
Gross et al. (2017) MoE of pretrained experts CNN MoE 2.2.2
Lin et al. (2017) Dynamic channel pruning CNN Channel 2.2.1
Odena et al. (2017) Combination of MoE and Early Exits CNN MoE 2.2.2
Liu and Deng (2018) Selective execution with self-defined topolog MLP Computational path 2.2.1
Rosenbaum et al. (2018) Self-organizing network for multi-modal learning CNN Computational path 2.2.1
Shazeer et al. (2018) MoE for features of visually similar classes CNN MoE 2.2.2
Veit and Belongie (2018) Network with adaptive inference graphs CNN Computational path 2.2.1
Wang et al. (2018) RNN for skip decision of CNN blocks CNN Block 2.2.1
Wang et al. (2018) Selectively skip Convolutional blocks based on the previous layer CNN Block 2.2.1
Wu et al. (2018) RL policy network for residual blocks skipping CNN Block 2.2.1
Chen et al. (2019) Global gating module for channel selection CNN Channel 2.2.1
Gao et al. (2019) Skip negligible input and output channels CNN Channel 2.2.1
Hua et al. (2019) Learn specialized convolutional kernels as combination of learned experts CNN MoE 2.2.2
Abbas and Andreopoulos (2020) MoE with inductive prior bias to certain experts CNN MoE 2.2.2
Feng et al. (2020) Models the relationship between the sample space and the latent routing space CNN Computational path 2.2.1
Jiang et al. (2020) MoE combined with layer skipping CNN MoE 2.2.2
Li et al. (2020) Adapts the scale at which each image gets processed CNN - 2.2.1
Pavlitskaya et al. (2020) MoE for interpretability of Semantic Segmentation CNN MoE 2.2.3
Shen et al. (2020) Layer-wise adaptive quantization and skip CNN Layer 2.2.1
Verelst and Tuytelaars (2020) Spatially execute convolutional filters only on important image patches CNN Channel 2.2.1
[…]
@@
Paper Summary Architecture Dynamic component Section
Wang et al. (2020) MoE at the level of convolutional filters, composed on-the-fly CNN MoE 2.2.2
Wang et al. (2020) Layer and channel skipping for IoT applications CNN Layer & channel 2.2.3
Colleman et al. (2021) Investigates hardware constraints in the context of DyNN CNN - 2.2.3
Hazimeh et al. (2021) Address the problem of slow convergence in sparse gates CNN MoE 2.2.2
Li et al. (2021) Highlight the informative features in both the channel and spatial dimensions CNN Channel 2.2.1
Park et al. (2021) Adaptive routing path for Object Detection CNN Computational path 2.2.3
Riquelme et al. (2021) MoE in the context of ViT Transformer MoE 2.2.2
Sun et al. (2021) Adaptive model precision and frame skipping for Video Recognition CNN - 2.2.3
Wang et al. (2021) Adaptive patch location for Video Recognition CNN Patch location 2.2.3
Liang et al. (2022) Method to accelerate MoE for multi-task ViT Transformer MoE 2.2.2
Liu et al. (2022) Self-regulate computations according to performance target and resources available CNN Computational path 2.2.1
Meng et al. (2022) Block skipping in Transformer Transformer Block 2.2.1
Wang et al. (2022) One stage adaptive patch location for Video Recognition CNN Patch location 2.2.3
Wang et al. (2022) Adaptive patch location for Video Recognition with conditional exits CNN Patch location 2.2.3
Gao et al. (2023) Spatial and channel pruning hardware accelerator CNN Channel 2.2.3
Pan et al. (2023) Studies adversarial attacks on dynamic models CNN - 2.2.3
Verelst and Tuytelaars (2023) Adjusts dynamically the processing resolution of image regions CNN Computational path 2.2.1
Zhang et al. (2023) Discuss adversarial attacks and robustness in the context of MoE CNN MoE 2.2.2
Devoto et al. (2024) Layer skipping in the fine-tuning of ViT Transformer Block 2.2.1
Liu et al. (2024) Studies different approaches of MoE in the context of ViT Transformer MoE 2.2.2
Yang et al. (2024) Adapts kernel weights and receptive field for different frames in Temporal Action Detection CNN Channel 2.2.3
[…]
@@ 2.3 Token Skimming
More formally, given a token matrix $X\in R^{N\times D}$, where $N$ is the number of tokens and $D$ is the dimension of each token, the skimming operation can be defined as a function $f:R^{N\times D}\rightarrow R^{N^{\prime}\times D}$, where $N^{\prime}\leq N$. The construction of the function $f$ can be implemented in different ways:
  • In Section 2.3.1 we discuss the major contributions in token dropping, where $f$ becomes a selection function on the subset of tokens that can either be dropped (hierarchical skimming) or considered for selection again in the next block (per-block skimming).
@@
Method Accuracy (top-1) GFLOPS Image Throughput Params (M) Type
DynamicViT-LV-S/0.7 (Rao et al., 2021) 83.0 4.6 - 26.9 Drop
S2ViTE-Small (Chen et al., 2021b) 79.2 - - 14.6 Drop
Evo-ViT (Xu et al., 2022) 79.4 - - - Merge
E-ViT-DeiT-S (Liang et al., 2022) 81.3 - 4385 - Merge
ATS (Fayyaz et al., 2022) 79.7 2.9 - 22.1 Drop
SaiT (Li et al., 2022) 83.1 4.3 - 26.2 Drop
A-ViT (Yin et al., 2022) 78.6 3.6 - 22 Drop
(Long et al., 2023) 79.6 3.0 - 22.1 Merge
ToMe (Bolya et al., 2023) 79.3 - 1564 - Merge
dTPS (Wei et al., 2023) 80.1 3.0 - 22.8 Merge
MSViT (Havtorn et al., 2023) 79.4 4.1 - - Token number
ToMeCIS (Seol et al., 2023) 78.9 2.7 1950 22.1 Merge
NEPAM (Jiang et al., 2023) 79.4 3.7 2452 - Merge
ToFu (Kim et al., 2024) 79.6 2.7 1561 - Drop & merge
GTP (Xu et al., 2024) 79.5 - 1581 22.1 Merge
IdleViT (Xu et al., 2024) 79.9 - 3031 22.1 Drop
@@ 2.3.1 Token dropping
Hierarchical skimming
In one of the early works on hierarchical skimming, Rao et al. (2021) introduces DynamicViT, with the idea of pruning redundant tokens progressively and dynamically. The sparsification of the tokens happens at predetermined layers within the network. During the training phase, a customized binary decision mask is produced. The purpose of the mask is to separate the informative and uninformative tokens. A local account of the token importance is achieved by reprojecting the tokens into a smaller latent space, while a global account is obtained by average pooling over the local importance. The two importance tokens are then concatenated, processed by a MLP, and reprojected through a softmax to obtain the probability for each original token. The Gumbel-Softmax trick is used to overcome the non-differentiable problem of sampling from a discrete distribution, while an attention-masking strategy guarantees that the excluded tokens are kept and can again be influential in the following operation down the model during the training process. This serves the purpose of keeping the correct batch size among all images during training, while in the inference phase, the unimportant tokens can be dropped to speed up the process. On the other hand, Pan et al. (2021) proposes a different method to hierarchically skim redundant tokens. It optimizes the network in a Curriculum Learning manner, training a router component for each block of heads and subsequently fixing it and moving on to the next block. The optimization of the routers is done through the REINFORCE method (Williams, 1992), where the reward function takes into account both efficiency and accuracy, and subsequently fine-tunes the Transformer blocks with a gradient computed based on the cross-entropy loss. The paper highlights how these types of methods are useful to provide an interpretability of the model behavior given the input. The model has been tested not only for classification but also for segmentation and video action recognition, showing improvement in all tasks. Li et al. (2022) proposes an adaptive pruning strategy that dynamically drops the number of tokens based on the weights of the attention block. In this case, a Token Importance Score (TIS) is calculated during training and inferred starting from the attention score at a given head. The tokens are then filtered based on a top-$k$ strategy considering a fixed number of tokens, or with a dynamic selection where a fixed threshold is set and only the tokens above the threshold pass. Knowledge distillation from a teacher model is also introduced to improve the ability of early layers to learn the TIS.
@@
Per-block skimming
Mersha (2022) introduces the same concept of block skipping for Transformer architectures, but on a token level instead of the entire image. This is possible due to the different nature of the Transformer architecture: an image is treated as a set of separate tokens that interact with each other only in the Self-Attention Block. Skipping blocks on a per-token case is conceptually equivalent to deciding which ones among all available tokens attend the Transformer block. This allows an oracle function to evaluate all the tokens before each Transformer block and decide which ones should be processed by the block and which ones should skip to the next block. In the presented paper, the generated score for each sample is stored and fed to the oracle of the successive layer, which allows the network to propagate these decisions up to the final layer, where the quality of the decision is estimated and summed to the final loss. Also, Xu et al. (2024) proposes a very similar method, where tokens are also selected to either attend the block or skip to the next block. In this case, the model uses the attention score towards the classification token to determine the amount of tokens that will pass to the block. Furthermore, it proposes an optimization problem based on the binary graph cut technique, where the optimization tries to find the minimum normalized cut of tokens.
A variation of this concept is proposed in Lei (2023), where each block passes through an MLP layer, which outputs a score. All the generated scores are then passed through a softmax, and top-$k$ are selected to attend the Tranformer block. The paper tests the solution on a large variety of tasks from very different domains.

### v3

@@ 2.2.2 Mixture of Experts
Wang et al. (2020a) presents an architecture based on MoE that composes the filter of a CNN on the fly from a set of (expert) channels. The input is fed into both a base Convolutional Network and a shallow embedding network. The embedding network outputs a latent mixture of weights, which is used by the convolutional network to select the expert channels to use for each specific layer. The paper also demonstrates that this gating method is equivalent to constructing a MoE for each output channel. A loss based on the computational cost is also used to encourage sparsity in the gating network. Also, Hua et al. (2019) presents a similar idea and introduces a revision of the MoE method to learn specialized convolutional kernels for each example. This increases the size and capacity of the network while maintaining efficient inference. The idea is to compute convolutional kernels as a linear combination of $n$ learned experts. In this way, the convolutional kernel is constructed on the fly based on the input needs.
@@
In Hazimeh et al. (2021) the authors raise the problem of slow convergence in the case of sparse gates (like top-$k$), which results in performance issues when training with gradient-based methods. The authors then introduce a reformulation of the binary selection based on first-order methods (such as SGD). Zhang et al. (2023a) presents a method to strengthen the MoE CNNs. The paper talks mostly about adversarial attacks and robustness, trying to understand what, among all the factors (network, router), is mostly responsible for the lack of robustness and proposing a method to train robust MoE CNNs.
Riquelme et al. (2021) is the first to introduce the use of MoE in the context of ViT models. The specialization in this case can be performed on a token level, where each token gets assigned to one (or more) experts, and experts can accept tokens up to a capacity level. Generally, in ViTs, the experts is applied to the MLP portion of the network. The paper also introduces a routing algorithm that repurposes model sparsity to skip the computation of some patches, reducing computations on uninformative image regions. Liu et al. (2024a) further studies different routing approaches of MoE in the context of ViTs. The comparison considers binary routers for hard assignment among experts, and soft routers for soft assignment between experts and weighted combinations of the expert output tokens. Further assignment divisions can be made, either matching experts to tokens or vice versa. The study arrives at the conclusion that the soft MoE outperforms the binary solution in terms of accuracy and thus is favorable. M3ViT (Liang et al., 2022a) introduces the use of a MoE framework to accelerate multi-task ViTs. Sparse task-specific experts get activated during training, which disentangles the parameter spaces to avoid different tasks’ training conflicts. The solution proposes tokens to per-expert queues to enable expert-by-expert computation rather than token-by-token. At inference time, this allows for activating only the task-corresponding sparse expert. A computation reordering scheme is also introduces to switch between tasks for deployment on memory-constrained devices.
In Jain et al. (2024), the idea of a nested experts structure is presented, where the model learns to dynamically process less important tokens with a less detailed representation. A novel routing algorithm for the tokens’ assignment to experts is also proposed. Xie et al. (2024) tries to overcome the limited amount of tokens that each expert sees by introducing a tailored mutual distillation among experts to allow each expert to get a better perception of the global task.
We refer the reader to Han et al. (2024) for an up-to-date, in-depth study on the design of MoE in the context of ViTs.
@@ 2.2.3 Applications
In this section, we go through publications that present the application of computational routing techniques to solve specific applicational problems. Park et al. (2021) explore the use of dynamic routing networks in the context of object detection. The proposed architecture is based on MSDNet (Huang et al., 2018) and automatically generates a sample-adaptive routing path on the fly. Two additional metrics for a global budget constraint and local path similarity regularization are added. The first serves the purpose of inducing sparsity in the network, while the second aims to generate more diverse routing paths. Liu et al. (2025) explores the use of MoE models to dynamically achieve low latency and high accuracy as needed in the context of 3D object detection. After multiscale preprocessing, the system dispatches the input to different experts based on object distance and visibility.
Pavlitskaya et al. (2020)
@@
investigates a MoE architecture to achieve interpretability in the context of semantic segmentation while retaining comparable result quality. The principle consists in being able to observe both the overall model output as well as individual expert outputs, where the agreement or disagreement between those partial solutions can be used for insights into the decision process.
@@
In Sun et al. (2021), a framework for efficient video recognition is presented. The solution selects the optimal precision for each frame conditioned on the input. Through Gumbel-Softmax sampling, an end-to-end differentiable approach selects optimal precision, taking both accuracy and efficiency into account. The solution is composed of a lightweight policy network (composed of a feature extractor and a LSTM module) and a video recognition network. The method also allows to skip frames in addition to selecting the precision dynamically. Furthermore, knowledge distillation from the full-precision recognition network is performed to guide the training of lower precisions and to improve accuracy. Yang et al. (2024a) presents DyFADet, a module that can adapt both kernel weights and receptive fields at different timestamps in the task of temporal action detection. The architectures dynamically adjust detector parameters when applied at different pyramid levels, which corresponds to detecting the actions with different time ranges. Sun et al. (2025) applies a MoE strategy to reduce the computational complexity of elaborating hyperspectral images in the task of object tracking.
@@
Furthermore, Wang et al. (2021a) introduces an adaptive patch localization solution for video recognition. By using a Reinforcement Learning approach, the patch localization problem is defined as a sequential decision task, and a CNN is used to quickly process the full video sequence. Then, the selected patches are used by a larger network to get the final prediction. The solution takes into account also the possibility of skipping some less valuable frames in the training policy. Wang et al. (2022a) reformulates AdaFocus as one-stage algorithm by introducing a differentiable interpolation-based patch selection approach. Furthermore, in this case, a conditional-exit technique is included to perform temporal adaptive computation. Wang et al. (2022b) improves on top of the second version of AdaFocus by activating the expensive high-capacity network only on some small but informative 3D video cubes. Huang et al. (2025) applies MoE to the point cloud registration task, introducing a prior-guided module to increase feature distinctiveness.
@@
Paper Summary Archit. Dynamic component Section Ahmed et al. (2016) Introduces MoE architecture for CNN CNN MoE 2.2.2 Bolukbasi et al. (2017) Ensemble of networks chained in an acyclic computation graph CNN Computational path 2.2.1 Figurnov et al. (2017) Dynamically adjusts the number of layers for certain regions of the image CNN Layer 2.2.1 Gross et al. (2017) MoE of pretrained experts CNN MoE 2.2.2 Lin et al. (2017) Dynamic channel pruning CNN Channel 2.2.1 Odena et al. (2017) Combination of MoE and Early Exits CNN MoE 2.2.2 Liu and Deng (2018) Selective execution with self-defined topolog MLP Computational path 2.2.1 Rosenbaum et al. (2018) Self-organizing network for multi-modal learning CNN Computational path 2.2.1 Shazeer et al. (2018) MoE for features of visually similar classes CNN MoE 2.2.2 Veit and Belongie (2019) Network with adaptive inference graphs CNN Computational path 2.2.1 Wang et al. (2018b) RNN for skip decision of CNN blocks CNN Block 2.2.1 Wang et al. (2018a) Selectively skip Convolutional blocks based on the previous layer CNN Block 2.2.1 Wu et al. (2018) RL policy network for residual blocks skipping CNN Block 2.2.1 Chen et al. (2019b) Global gating module for channel selection CNN Channel 2.2.1 Gao et al. Skip negligible input and output channels CNN Channel 2.2.1 Hua et al. (2019) Learn specialized convolutional kernels as combination of learned experts CNN MoE 2.2.2 Abbas and Andreopoulos (2020) MoE with inductive prior bias to certain experts CNN MoE 2.2.2 Feng et al. (2020) Models the relationship between the sample space and the latent routing space CNN Computational path 2.2.1 Jiang et al. (2020) MoE combined with layer skipping CNN MoE 2.2.2 Li et al. (2020b) Adapts the scale at which each image gets processed CNN - 2.2.1 Pavlitskaya et al. (2020) MoE for interpretability of Semantic Segmentation CNN MoE 2.2.3 Shen et al. (2020) Layer-wise adaptive quantization and skip CNN Layer 2.2.1 Verelst and Tuytelaars (2020) Spatially execute convolutional filters only on important image patches CNN Channel 2.2.1 Wang et al. (2020a) MoE at the level of convolutional filters, composed on-the-fly CNN MoE 2.2.2 Wang et al. (2020b) Layer and channel skipping for IoT applications CNN Layer & channel 2.2.3 Colleman et al. (2021) Investigates hardware constraints in the context of DyNN CNN - 2.2.3
@@
Paper Summary Archit. Dynamic component Section Hazimeh et al. (2021) Address the problem of slow convergence in sparse gates CNN MoE 2.2.2 Li et al. (2021) Highlight the informative features in both the channel and spatial dimensions CNN Channel 2.2.1 Park et al. (2021) Adaptive routing path for Object Detection CNN Computational path 2.2.3 Riquelme et al. (2021) MoE in the context of ViT Transformer MoE 2.2.2 Sun et al. (2021) Adaptive model precision and frame skipping for Video Recognition CNN - 2.2.3 Wang et al. (2021a) Adaptive patch location for Video Recognition CNN Patch location 2.2.3 Liang et al. (2022a) Method to accelerate MoE for multi-task ViT Transformer MoE 2.2.2 Liu et al. (2022) Self-regulate computations according to performance target and resources available CNN Computational path 2.2.1 Meng et al. (2022) Block skipping in Transformer Transformer Block 2.2.1 Wang et al. (2022a) One stage adaptive patch location for Video Recognition CNN Patch location 2.2.3 Wang et al. (2022b) Adaptive patch location for Video Recognition with conditional exits CNN Patch location 2.2.3 Gao et al. (2023) Spatial and channel pruning hardware accelerator CNN Channel 2.2.3 Pan et al. (2023) Studies adversarial attacks on dynamic models CNN - 2.2.3 Verelst and Tuytelaars (2023) Adjusts dynamically the processing resolution of image regions CNN Computational path 2.2.1 Zhang et al. (2023a) Discuss adversarial attacks and robustness in the context of MoE CNN MoE 2.2.2 Devoto et al. (2025) Layer skipping in the fine-tuning of ViT Transformer Block 2.2.1 Liu et al. (2024a) Studies different approaches of MoE in the context of ViT Transformer MoE 2.2.2 Yang et al. (2024a) Adapts kernel weights and receptive field for different frames in Temporal Action Detection CNN Channel 2.2.3 Han et al. (2024) Empirical study on MoE design Transformer MoE 2.2.2 Jain et al. (2024) The model learns to dynamically process less important tokens with a less detailed representation Transformer MoE 2.2.2 Xie et al. (2024) Distillation among experts to allow each expert to get a better perception Transformer MoE 2.2.2 Huang et al. (2025) MoE applied to Point Cloud registration task Transformer MoE 2.2.3 Liu et al. (2025) Multiscale MoE for low latency and high accuracy in 3D object detection Transformer MoE 2.2.3 Sun et al. (2025) MoE strategy for reducing hyperspectral images computational complexity Transformer MoE 2.2.3
@@ 2.3 Token Skimming
We use token skimming as an umbrella term to denote two distinct token reduction strategies: token dropping and token merging. More formally, given a token matrix $X\in R^{N\times D}$, where $N$ is the number of tokens and $D$ is the dimension of each token, the skimming operation can be defined as a function $f:R^{N\times D}\rightarrow R^{N^{\prime}\times D}$, where $N^{\prime}\leq N$. The construction of the function $f$ can be implemented in different ways:
  • 1.
    In Section 2.3.1 we discuss the major contributions in token dropping, where $f$ becomes a selection function on the subset of tokens that can either be dropped (hierarchical dropping) or considered for selection again in the next block (per-block dropping).
  • 2.
@@
Method Accuracy (top-1) GFLOPS Image Throughput Params (M) Type DynamicViT-LV-S/0.7 (Rao et al., 2021) 83.0 4.6 - 26.9 Drop S2ViTE-Small (Chen et al., 2021b) 79.2 - - 14.6 Drop Evo-ViT (Xu et al., 2022) 79.4 - - - Merge E-ViT-DeiT-S (Liang et al., 2022b) 81.3 - 4385 - Merge ATS (Fayyaz et al., 2022) 79.7 2.9 - 22.1 Drop SaiT (Li et al., 2022) 83.1 4.3 - 26.2 Drop A-ViT (Yin et al., 2022) 78.6 3.6 - 22 Drop (Long et al., 2023) 79.6 3.0 - 22.1 Merge ToMe (Bolya et al., 2023) 79.3 - 1564 - Merge dTPS (Wei et al., 2023) 80.1 3.0 - 22.8 Merge MSViT (Havtorn et al., 2023) 79.4 4.1 - - Token number ToMeCIS (Seol et al., 2023) 78.9 2.7 1950 22.1 Merge NEPAM (Jiang et al., 2023) 79.4 3.7 2452 - Merge ToFu (Kim et al., 2024) 79.6 2.7 1561 - Drop & merge GTP (Xu et al., 2024c) 79.5 - 1581 22.1 Merge IdleViT (Xu et al., 2024b) 79.9 - 3031 22.1 Drop Soft Token Merging Yuan et al. (2024) 79.3 2.3 4566 24.0 Merge
@@ 2.3.1 Token dropping
Hierarchical dropping
In one of the early works on hierarchical dropping, Rao et al. (2021) introduces DynamicViT, with the idea of pruning redundant tokens progressively and dynamically. The sparsification of the tokens happens at predetermined layers within the network. During the training phase, a customized binary decision mask is produced. The purpose of the mask is to separate the informative and uninformative tokens. A local account of the token importance is achieved by reprojecting the tokens into a smaller latent space, while a global account is obtained by average pooling over the local importance. The two importance tokens are then concatenated, processed by a MLP, and reprojected through a softmax to obtain the probability for each original token. The Gumbel-Softmax trick is used to overcome the non-differentiable problem of sampling from a discrete distribution, while an attention-masking strategy guarantees that the excluded tokens are kept and can again be influential in the following operation down the model during the training process. This serves the purpose of keeping the correct batch size among all images during training, while in the inference phase, the unimportant tokens can be dropped to speed up the process. On the other hand, Pan et al. (2021) proposes a different method to hierarchically skim redundant tokens. It optimizes the network in a Curriculum Learning manner, training a router component for each block of heads and subsequently fixing it and moving on to the next block. The optimization of the routers is done through the REINFORCE method (Williams, 1992), where the reward function takes into account both efficiency and accuracy, and subsequently fine-tunes the Transformer blocks with a gradient computed based on the cross-entropy loss. The paper highlights how these types of methods are useful to provide an interpretability of the model behavior given the input. The model has been tested not only for classification but also for segmentation and video action recognition, showing improvement in all tasks. Li et al. (2022) proposes an adaptive pruning strategy that dynamically drops the number of tokens based on the weights of the attention block. In this case, a Token Importance Score (TIS) is calculated during training and inferred starting from the attention score at a given head. The tokens are then filtered based on a top-$k$ strategy considering a fixed number of tokens, or with a dynamic selection where a fixed threshold is set and only the tokens above the threshold pass. Knowledge distillation from a teacher model is also introduced to improve the ability of early layers to learn the TIS.
@@
Per-block dropping
Mersha (2022) introduces the same concept of block skipping for Transformer architectures, but on a token level instead of the entire image. This is possible due to the different nature of the Transformer architecture: an image is treated as a set of separate tokens that interact with each other only in the Self-Attention Block. Skipping blocks on a per-token case is conceptually equivalent to deciding which ones among all available tokens attend the Transformer block. This allows an oracle function to evaluate all the tokens before each Transformer block and decide which ones should be processed by the block and which ones should skip to the next block. In the presented paper, the generated score for each sample is stored and fed to the oracle of the successive layer, which allows the network to propagate these decisions up to the final layer, where the quality of the decision is estimated and summed to the final loss. Also, Xu et al. (2024b) proposes a very similar method, where tokens are also selected to either attend the block or skip to the next block. In this case, the model uses the attention score towards the classification token to determine the amount of tokens that will pass to the block. Furthermore, it proposes an optimization problem based on the binary graph cut technique, where the optimization tries to find the minimum normalized cut of tokens.
A variation of this concept is proposed in Lei et al. , where each block passes through an MLP layer, which outputs a score. All the generated scores are then passed through a softmax, and top-$k$ are selected to attend the Tranformer block. The paper tests the solution on a large variety of tasks from very different domains.